# Privacy for Quantum Annealing. Attack on Spin Reversal Transformations in the case of cryptanalysis

### v1

@@ Abstract
This paper demonstrates that applying spin reversal transformations (SRT), commonly known as a sufficient method for privacy enhancing in problems solved using quantum annealing, does not guarantee privacy for all possible problems. We show how to recover the original problem from the Ising problem obtained using SRT when the resulting problem in Ising form represents the algebraic attack on the $E_{0}$ stream cipher. A small example is used to illustrate how to retrieve the original problem from the one transformed by SRT. Moreover, it is shown that our method is efficient even for full-scale problems.
@@ 1 Introduction
Quantum annealing is a highly complex process. In this paper, we focus only on its applications and the confidentiality of the data used in the calculations. In the following section, the aspects of the communication scheme assumed, and the calculations used are discussed.
1.1 Quantum cloud services and privacy.
Quantum cloud services are now becoming more and more popular. The delivered products allow us to solve a wide range of problems. In this paper, we focus on two specific quantum services:
@@
As presented in [3], quantum annealing has many applications. It can be used in areas such as:
  • traffic flow optimization;
  • logistic, vehicle routing problem;
  • in finance, portfolio optimization;
  • quantum simulation.
However, typically, solutions are implemented using remote access. Such a solution requires the transfer of data to an outsourcer. Under ideal conditions, such a scheme can be depicted as in the figure 1.
Figure 1: Scheme of ideal communication.
The tasks to be solved in both cases require representation using the Ising model. In this case, communication with the quantum cloud requires sending a connection matrix $J$ and a vector of biases $h$, presented as equations successively (3) and (2). The communication scheme is straightforward, as presented in the figure 1. Client prepares problem in Ising model (in the case of Quantum Annealing or Quantum Approximate Optimization Algorithm) and then sends it to Quantum Cloud Service via the internet. After computations are done, the solution is sent to the client similarly.
In specific cases, such an arrangement is unacceptable. As presented in [4], [5] or [6] quantum annealing can be used in cryptoanalysis. In such use, it is not permissible to transfer the data to an external party. The answer is privacy-preserving methods.
Privacy in quantum computing is a promising area of research. Secure protection algorithms will expand the market for quantum services. As it said in [7], Secure Quantum Computing can be divided into two groups of methods:
  • Blind Quantum Computing is a technique that allows outsourcing computations without disclosing the computation’s details to the server. This group of protocols is often impractical and requires the customer to have the quantum hardware or the participation of several servers.
  • Quantum Homomorphic is encryption that involves performing calculations on encrypted data. The result of the calculation, after decryption, is the result of the original problem.
This paper focuses on a method called spin reversal transformation (SRT) presented in [8]. The authors propose homomorphic encryption for quantum annealing, which they believe protects the details of quantum annealing instances against a malicious cloud. As said in the paper, the scheme runs on the Ising model, so it can also be used for privacy preservation in the Quantum Approximate Optimization Algorithm.
1.2 Man in the middle. The real threat.
@@
The communication diagram shown in Figure 1 has little to do with actual communication conditions. With the growing number of cloud service providers, the threat is growing. Providers services are highly needed, and their availability is severely limited. This leads to a situation where a new provider offering competitive services can gain popularity rapidly. The internal structure of cloud services remains a mystery, and it is not known if there is no eavesdropping between the client and the actual computer. Such a new, untrusted provider may be malicious, and communications between the user and the actual quantum computer may be intercepted.
Figure 2 shows a more realistic scheme. Between client and quantum cloud service is the adversary. All communication goes through the adversary, which has any possibility to perform operations on the received problem. This paper assumes that adversaries eavesdrop on intercepted communications and pass them forward. Intercepted data can lead to financial losses (for commercial applications) or even threaten security (in the case of military applications such as cryptanalysis).
Figure 2: Communication scheme with a potential adversary.
@@ 1.3 A description of the models used
As mentioned earlier, using the Quantum Approximate Optimization Algorithm or Quantum Annealing requires presenting the problem in an sepcific form. One such form is the Ising model, known since 1920. Ernst Ising and Wilhelm Lenz introduced it as a description of magnetic materials. Despite its original application, the model has also been applied in combinatorial optimization. As presented in [9] there exist Ising formulations of many NP problems such as:
  • Graph partitioning [10];
  • Knapsack problem [11];
  • Graph coloring [11];
  • Traveling salesman [11].
Moreover, the model can be applied in cryptanalysis. Including but not limited to the factorization presented by Jiang, in [4], breaking the discrete logarithm [12, 13, 14, 15, 16] or attacking symmetric ciphers [5, 17, 6]. It is worth noting that the last two applications require a transition between the QUBO model and the Ising model.
As mentioned in [18], each Ising problem can be viewed as a minimizing expression presented as Equation (1).
$f(s)=\sum_{i=0}^{n-1}h_{i}s_{i}+\sum_{i,j=0}^{n-1}J_{i,j}s_{i}s_{j}$ (1)
Vector $s$ is called state and each variable $s_{i}\in\{-1,1\}$ is called spin. In practice, the following are used to characterize the problem:
@@
An alternative model is the QUBO model, which uses binary variables. Quadratic unconstrained binary optimization can also be viewed as minimizing specific expression, shown as Equation (4).
$f(x)=\sum_{i=0}^{n-1}Q_{i,i}x_{i}+\sum_{i,j=0}^{n}Q_{i,j}x_{i}x_{j}.$ (4)
Vector $x$ contains $n$ binary variables.
As mentioned earlier, there is a simple transformation between the Ising model and the QUBO problem. Its main idea is shown in the Figure 3. The transition between models requires simple substitutions:
  • from Ising model to QUBO:
    $x_{i}=\frac{1}{2}\cdot(s_{i}+1),$ (5)
  • from QUBO to Ising model:
    $s_{i}=2x_{i}-1.$ (6)
In the rest of the paper, we skip transitions between models, considering the operation trivial.
Figure 3: Transformation between models.
@@
$\begin{array}[]{c}Q_{i,j}=4J_{i,j},\\ Q_{i,i}=2(h_{i}-\sum_{j}J_{i,j}-\sum_{j}J_{j,i}.)\end{array}$ (7)
@@ 2 Spin Reversal Transformation in detail
The simplest way to protect privacy is to encrypt the problem using a homomorphic algorithm. The first ideas of privacy protection in this way were described in [19]. However, they were not very popular for a long time. The idea of applying such a method to quantum cloud computing is described in [8] and partly in [7]. The method utilizes a random sequence to reverse a given sign in an Ising problem instance. It is important to note that the minimal energy of the instance does not change. As depicted in [8] and presented using the Equation (12), the solution to the concealed problem, when uncovered, is the solution to the original problem. In addition, the authors of the paper demonstrate that their method is efficient.
2.1 Secure communication scheme
The idea of a secure scheme is presented in Figure 4. The communication channel is secured by encrypting the problem being sent. As presented in [8], a suitable algorithm will recover the solution to the original problem from the encrypted solution. After preparing the problem, the client will perform encryption and send the new problem to the Quantum Cloud Service. An adversary, having intercepted such a problem, according to the authors, will not be able to reconstruct the original problem without knowing the original key. After processing, the solution to the encrypted problem is sent to the client. Using a previously used key, the customer can recover the solution to the original problem.
Figure 4: A secure communication scheme, proposed in [8].
@@ 2.2 Description of the algorithm
The encryption scheme described in [8] is based on SRT, called gauge transformation [20]. The transformation uses a binary string $x$ to change the signs of the selected coefficients. After applying mentioned transformation, Equation (1) is transformed as follows:
@@
The same sequence $x$ must be used to determine the original solution. The corresponding solution can be determined according to the Equation 11.
$s_{i}=(-1)^{x_{i}}s_{i}^{*}$ (11)
As mentioned in [8] and [20], this transformation does not change the ground state. Furthermore, a solution $s$ minimizes $f(s)$ if and only if solution $s^{*}$ minimizes $f^{*}(s^{*})$. This follows from the equality shown in the Equation below (12)
$\begin{array}[]{c}f^{*}(s^{*})=\sum_{i=0}^{n-1}(-1)^{x_{i}}h_{i}\cdot(-1)^{x_{i}}s_{i}+\sum_{i,j=0}^{n-1}(-1)^{x_{i}+x_{j}}J_{i,j}\cdot(-1)^{x_{i}}s_{i}\cdot(-1)^{x_{j}}s_{j}=\\ \sum_{i=0}^{n-1}(-1)^{2x_{i}}h_{i}s_{i}+\sum_{i,j=0}^{n-1}(-1)^{2x_{i}+2x_{j}}J_{i,j}s_{i}s_{j}=\\ \sum_{i=0}^{n-1}h_{i}s_{i}+\sum_{i,j=0}^{n-1}J_{i,j}s_{i}s_{j}=f(s)\\ \end{array}$ (12)
@@ 3 Details of the proposed attack
This paper argues that the privacy-enhancing mechanism presented in [8] fails in many instances. As a counterexample, we demonstrate that the method fails when the method of transforming the stream cipher cryptanalytic problem is known. Specifically, we show that, without prior knowledge of the random sequence used to conceal the given Ising problem for cryptanalysis of the $E_{0}$ cipher, one can easily retrieve the original problem using the intercepted data. Furthermore, by knowing the solution to the concealed problem, one can also deduce the solution to the original problem. It also allows the recovery of the concealment key and the data on which the client cares about confidentiality (key hidden in the solution to the problem sent, ciphertexts, or keystream). With a concealment key, the adversary can recover communications encrypted with this key.
@@ 3.1 The problem for the $E_{0}$ cipher
We will present an attack on the mentioned scheme using the QUBO problem corresponding to the cryptanalysis of the $E_{0}$ cipher, presented in the [21]. As the authors point out, the reduction presented in the paper in the next few years can be practically embedded in a commercially available quantum annealer. This opens up the possibility of application to real cryptanalysis, for which it is essential to maintain the privacy of transmitted data and recovered keys.
The cipher is built from three elements:
  • 4 shift registers with linear feedback, specified by the following primitive polynomials $f_{i}(x)$:
    $\begin{array}[]{cc}L_{1}:&f_{1}(x)=x^{25}\oplus x^{20}\oplus x^{12}\oplus x^{8}\oplus 1,\\ L_{2}:&f_{2}(x)=x^{31}\oplus x^{24}\oplus x^{16}\oplus x^{12}\oplus 1,\\ L_{3}:&f_{3}(x)=x^{33}\oplus x^{28}\oplus x^{24}\oplus x^{4}\oplus 1,\\ L_{4}:&f_{4}(x)=x^{39}\oplus x^{36}\oplus x^{28}\oplus x^{4}\oplus 1.\end{array}$ (15)
@@
    $F_{1}:s_{t+1}=\Bigl{\lfloor}\frac{\sum_{i=1}^{4}x_{i}+c_{t}}{2}\Bigr{\rfloor}$
  • Blend Register, calculating the two-bit value of $c_{t+1}$, using the bijections described in the cipher specification:
    $F_{2}:c_{t+1}=s_{t+1}\oplus T_{1}[c_{t}]\oplus T_{2}[c_{t-1}].$
@@
Each bit of the key stream is described by 8 equations:
@@ 3.2 How to identify the problem and parameterize it?
Quantum service providers require many data to identify the user. Among such data, we can distinguish:
@@
With their use, an untrusted provider can identify the client and the research area in which the client works. Identifying the area of study will allow the data obtained to be matched with one of the known problems. Problems vary significantly in size, density, and coefficient range. Betting that resources are limited and methods cannot change the connectivity (density) or size of the problem.
The critical element of the attack is determining a parameterized matrix for the problems under consideration. It should be noted that such a matrix is determined only once for a given problem. Determination of the matrix acts as a precomputation. The interactive attack phase uses pre-made parameterized matrices.
The details of the parameterization will depend on the type of cipher. The general principle remains the same. The construction for the $E_{0}$ cipher is discussed below. As shown in Equation (16), only one of the equations depends on the bits of the keystream. According to a method of transformation algebraic attack presented in [5], for each keystream bit, the component of the final polynomial will be computed as below:
$z_{t}+l_{t+1}+m_{t+7}+n_{t+1}+o_{t+7}+c^{0}_{t}-2K=0\Big{/}^{2}$ (17)
Finally, six variables depending on keystream bits are obtained: $2z_{t}\cdot l_{t+1},2z_{t}\cdot m_{t+7},2z_{t}\cdot n_{t+1},2z_{t}\cdot o_{t+7},2z_{t}\cdot c_{t}^{0},-4z_{t}\cdot K$, note that K, is an integer variable and should be replaced with a binary variable. The coefficients resulting from the squaring of other equations are constant. According to Equations (8), equations dependent on keystream occurring in the vector of biases can be computed.
The alternative method does not require the direct analysis of equations. As shown above, each of the coefficients depends on the keystream bits linearly. A similar relationship will be satisfied for other stream ciphers.
Assume that an oracle $\phi$ is given which, for any keystream, will return an Ising model, as Equation (19). The oracle can be constructed based on known publications on selected attacks. A detailed description of the transformation of the algebraic attack to the QUBO problem, as in [5], [6], [21], allows the construction of the assumed oracle. In known cases, the construction of such an oracle does not require a large number of resources. As in the method requiring analysis of each equation, the oracle must be queried only in preparing parameterized matrices for attacks.
The parameterized matrix can be determined with such knowledge, as discussed below. Keystream with a hamming weight of 1 will be used to determine the matrix. Such stream can be demoted as $s_{i}$ and is presented as Equation (18).
@@
We denote $h_{i}$, $J_{i}$ as Ising model, such that:
@@
Additionally, we denoted $h_{\infty}$, $J_{\infty}$ as the Ising model for a stream with Hamming weight equals zero. Then, the parametrized problem is computed as follows:
$\begin{array}[]{c}h_{P}=h_{\infty}+\sum_{i=0}^{n}(h_{i}-h_{\infty})\cdot z_{i},\\ J_{P}=J_{\infty}+\sum_{i=0}^{n}(J_{i}-J_{\infty})\cdot z_{i},\end{array}$ (20)
where, $z_{i}$ is variable.

### v3

@@ Abstract
This paper demonstrates that applying spin reversal transformations, commonly known as a sufficient method for privacy enhancement in problems solved using quantum annealing, does not guarantee privacy for all possible cases. We show how to recover the original problem from the Ising problem obtained using spin reversal transformation when the resulting problem in Ising form represents the algebraic attack on the $E_{0}$ stream cipher. A small example illustrates how to retrieve the original problem from that transformed by spin reversal transformation. Moreover, we show that our method is efficient also for full-scale problems.
@@ 1 Introduction
Quantum optimization is a highly complex process. Despite this, it is gaining considerable popularity [1]. Due to its significant structural requirements, it is usually implemented as a cloud service. The delivered products allow us to solve a wide range of problems. In this paper, we focus on two specific aspects of quantum computation:
@@
As presented in [4, 5], among the application areas of annealing, we can distinguish traffic flow optimization, logistics, vehicle routing problems, finance, and quantum simulation. As interest in quantum optimization grows, so does the need for methods to keep the computations performed private. The classical homomorphic encryption approach [6, 7, 8] does not apply to quantum optimization. In this case, dedicated methods, such as [9], must be used.
In this paper, we focus only on the application method described in [9] and the confidentiality of the data used in the calculations. We attack the scheme shown in [9]. A pen-and-paper example of the attack supports the theoretical description of the attack.
This paper is organized as follows. Section 2 provides an overview of possible privacy-preserving methods, the assumed flow of communication, and a description of the models used. Section 3 describes the attack framework and the attack method’s details. Section 4 introduces our attack, while Section 5 demonstrates its practical feasibility using a pen-and-paper example. Finally, Section 6 concludes the paper.
2 Background
This section discusses privacy aspects in communication flow and the assumptions made about the adversary. It also presents the basics of optimization models and their transformations.
2.1 Quantum cloud services and privacy
Figure 1 considers the communication flow. The client has a specific task that he wants to realize using optimization. In the first step, an optimization problem is formulated for the given task. The following optional step is to encrypt the resulting optimization problem. Assuming ideal communication conditions or having your own quantum computer, this step can be skipped as potentially unnecessary. However, in real-world conditions, failing to do this seriously threatens the privacy of the computations performed. After encryption, the problem is passed to a computer that can perform the indicated optimization. After optimization, a solution is returned. If the problem has been encrypted, it will be necessary to decrypt the solution received further.
Figure 1: Data flow considered.
Due to the high cost of quantum infrastructure solutions [10], quantum computing is typically implemented using remote access. Such a solution requires the transfer of data to an outsourcer. Due to the inability to verify the cloud service provider’s intentions and the lack of control over the data flow on its side, the model discussed below may be assumed.
@@ 2.1 Quantum cloud services and privacy
The adversary is between the client and the quantum cloud service. All communication goes through the adversary, which has the ability to perform any operation on the received problem. This paper assumes that adversaries eavesdrop on intercepted communications and pass them forward. With the growing number of cloud service providers, the threat is increasing. Service providers are in high demand, and their availability is severely limited. This leads to a situation where a new provider offering competitive services can gain popularity rapidly. The internal structure of cloud services remains a mystery, and it is unknown if there is no eavesdropping between the client and the actual computer. Such a new, untrusted provider may be malicious, and communications between the user and the quantum computer may be intercepted.
In specific cases, such an arrangement is unacceptable. There are few critical applications of quantum optimization where the potential leak of private data involves serious consequences. Transferring the data to an external party is not permissible for such use. In the case of portfolio optimization, such a situation may lead to a financial benefit for the cloud service provider, which could use the optimal solution before passing it to the customer. Another branch of applications for which computational privacy is significant is cryptanalysis. As presented in [11], it may be used in factorizing numbers, solving the discrete logarithm [12, 13, 14, 15, 16] or [17, 18] in algebraic attacks on symmetric ciphers. Computing without privacy in this situation could compromise classified or strategic data.
Just as in the case of classical cloud computing, the solution to the problem is homomorphic encryption [8]; in the case of the discussed issue, the solution may be analogous to specialized privacy-preserving methods. Privacy in quantum computing is a promising area of research. Secure protection algorithms will expand the market for quantum services. As stated in [19], Secure Quantum Computing can be divided into two groups of methods:
  • Blind Quantum Computing is a technique that allows outsourcing computations without disclosing the computation’s details to the server. This group of protocols is often impractical and requires the customer to have quantum hardware or the participation of several servers.
  • Quantum Homomorphic Encryption involves performing calculations on encrypted data. The result of the calculation, after decryption, is the result of the original problem.
This paper focuses on the spin reversal transformation method presented mainly in [9]. The authors propose homomorphic encryption for quantum annealing, which they believe protects the details of quantum annealing instances against a malicious cloud. As stated in the paper, the scheme runs on the Ising model, so it can also be used for privacy preservation in the Quantum Approximate Optimization Algorithm.
@@ 2.2 A description of the models used
As mentioned earlier, using the Quantum Approximate Optimization Algorithm or Quantum Annealing requires presenting the problem in a specific form. One such form is the Ising model, which has existed since 1920. Ernst Ising and Wilhelm Lenz introduced it as a description of magnetic materials. Despite its original application, the model has also been applied in combinatorial optimization. As mentioned in [20], each Ising problem can be viewed as a minimizing expression presented as Equation (1)
$f(s)=\sum_{i=0}^{n-1}h_{i}s_{i}+\sum_{i,j=0}^{n-1}J_{i,j}s_{i}s_{j}.$ (1)
Vector $s$ is called the state, and each variable $s_{i}\in\{-1,1\}$ is called a spin. In practice, the following are used to characterize the problem:
@@
In this case, communication with the quantum cloud requires sending vector of biases $h$ and a connection matrix $J$, presented as Equations (2) and (3) respectively.
An alternative model is the QUBO model, which uses binary variables $x\in\{0,1\}$. Quadratic unconstrained binary optimization can also be viewed as minimizing the specific expression shown as Equation (4)
$f(x)=\sum_{i=0}^{n-1}Q_{i,i}x_{i}+\sum_{i,j=0}^{n}Q_{i,j}x_{i}x_{j},$ (4)
where vector $x$ contains $n$ binary variables.
As presented in [21], there exist Ising formulations of many NP problems, such as graph partitioning [22], the knapsack problem [23], graph coloring [23], or the traveling salesman [23] problem. However, it is easier for some tasks to formulate the problem in QUBO form. Such tasks include algebraic attacks on symmetric ciphers, as described in [17, 24, 18]. Applying privacy-preserving methods for such problems may require transitioning between the QUBO and Ising models and vice versa.
There is a simple transformation between the Ising model and the QUBO problem. Its main idea is to perform the following substitutions to change spin variables into binary variables. When moving from the Ising model to the QUBO:
$x_{i}=\frac{1}{2}\cdot(s_{i}+1).$
When moving from the QUBO to the Ising model:
$s_{i}=2x_{i}-1.$
@@
$\begin{array}[]{c}Q_{i,j}=4J_{i,j},\\ Q_{i,i}=2\bigl{(}h_{i}-\sum_{j}J_{i,j}-\sum_{j}J_{j,i}\bigr{)}.\end{array}$ (5)
@@
In the rest of the paper, we skip transitions between models, considering this operation trivial.
@@ 3 Spin Reversal Transformation in detail
This section briefly describes the application of spin reversal transformation to preserve privacy. This idea is mainly presented in [9] and partly in [19]. The method utilizes a random sequence to reverse a given sign in an Ising problem instance. According to the authors, an adversary, having intercepted a concealed problem, cannot reconstruct the original problem without knowing the original key.
@@ 3.1 Description of the algorithm
The encryption scheme described in [9] is based on spin reversal transformation, also called a gauge transformation [25]. The transformation uses a binary string $x$ to change the signs of the selected coefficients. After applying the mentioned transformation, Equation (1) is transformed as follows:
@@
The same sequence $x$ must be used to determine the original solution. The corresponding solution can be determined according to Equation 9:
$s_{i}=(-1)^{x_{i}}s_{i}^{*}.$ (9)
It is important to note that the minimal energy of the instance does not change. As shown in [9] and [25], the solution to the concealed problem, when uncovered, is the solution to the original problem.
@@ 4 Details of the proposed attack
This paper argues that the privacy-enhancing mechanism presented in [9] fails in many instances. As a counterexample, we demonstrate that the method fails when the technique of transforming the stream cipher cryptanalytic problem is known. Specifically, we show that, without prior knowledge of the random sequence used to conceal the given Ising problem for cryptanalysis of the $E_{0}$ cipher, one can easily retrieve the original problem using the intercepted data. Furthermore, by knowing the solution to the concealed problem, one can also deduce the solution to the original problem. It also allows the recovery of the concealment key and the data on which the client wants confidentiality (the key hidden in the solution to the problem sent, ciphertexts, or keystream). With the concealment key, the adversary can recover communications encrypted with this key. The design of our attack involves three phases:
  1. 1.
    Parameterization (optional): Based on knowledge of the type of optimization task and access to the oracle $\phi$, a parameterized matrix is constructed. It can be implemented before the attack and only once for a given optimization task. Subsequent attacks use the predetermined matrix.
  2. 2.
    System setup: Using the parameterized matrix and the intercepted encrypted optimization problem $h_{i}^{*},J_{i,j}^{*}$, a system of linear equations is created.
  3. 3.
    Solving: The obtained system of equations is solved, determining the keystream used to create the optimization task and the concealment key.
@@ 4.1 The problem for the $E_{0}$ cipher
We present an attack on the mentioned scheme using the QUBO problem corresponding to the cryptanalysis of the $E_{0}$ cipher, presented in [26]. As the authors point out, the reduction shown in that paper in the next few years can be practically embedded in a commercially available quantum annealer. This opens up the possibility of application to real cryptanalysis, for which it is essential to maintain the privacy of transmitted data and recovered keys.
Our proposed attack can also be applied to other highly structured optimization problems. The following example was chosen for its potential practicality and its scalability.
Below is a brief overview of the construction of the $E_{0}$ cipher. A full description of the cipher can be found in [27]. The cipher is built from three elements:
  • Four shift registers with linear feedback, specified by the following primitive polynomials $f_{i}(x)$:
    $\begin{array}[]{cc}L_{1}:&f_{1}(x)=x^{25}\oplus x^{20}\oplus x^{12}\oplus x^{8}\oplus 1,\\ L_{2}:&f_{2}(x)=x^{31}\oplus x^{24}\oplus x^{16}\oplus x^{12}\oplus 1,\\ L_{3}:&f_{3}(x)=x^{33}\oplus x^{28}\oplus x^{24}\oplus x^{4}\oplus 1,\\ L_{4}:&f_{4}(x)=x^{39}\oplus x^{36}\oplus x^{28}\oplus x^{4}\oplus 1.\end{array}$
@@
    $F_{1}:s_{t+1}=\Bigl{\lfloor}\frac{\sum_{i=1}^{4}x_{i}+c_{t}}{2}\Bigr{\rfloor}.$
  • Blend Register, calculating the two-bit value of $c_{t+1}$ using the bijections described in the cipher specification:
    $F_{2}:c_{t+1}=s_{t+1}\oplus T_{1}[c_{t}]\oplus T_{2}[c_{t-1}],$
    where $T_{i}$ are linear mappings:
    $\begin{array}[]{cl}T_{1}:&(x_{1},x_{0})\to(x_{1},x_{0}),\\ T_{2}:&(x_{1},x_{0})\to(x_{0},x_{1}\oplus x_{0}).\end{array}$
@@
Each bit of the keystream is described by eight equations:
@@
4.2 From algebraic attack to quantum optimization
This section briefly describes the transformation of the algebraic attack to the QUBO optimization problem. The main idea was presented in [17] and adapted to $E_{0}$ in [26]. In the first step, a system of equations is generated. To recover the initial internal state of the cipher, we need 128 bits of the keystream. As mentioned, each stream bit is described by eight equations. In total, the system describing the cipher will have 1024 equations. Below, we show how to perform transformations of the resulting system in a few steps:
  1. 1.
    Equations $f_{i},i=\overline{0,6}$ are transformed into equations with binary variables and integer coefficients. Equation $f_{7}$ does not require such a transformation. From the construction of the cipher, it is already in this form:
    $f_{i}^{\prime}\equiv 0(\text{mod }2)\to f_{i}-2k_{i}=0.$
    Each integer variable $k_{i}$ is bounded, $k_{i}\leq\lfloor\frac{f_{i}^{\text{max}}}{2}\rfloor$, where $f_{i}^{\text{max}}$ is the maximum value of the selected polynomial.
  2. 2.
    In the standard transformation, the equations must be linearized in the next step. However, this step is skipped in this paper; the $E_{0}$ design induces no nonlinear equations.
  3. 3.
    Then variables $k_{i}$ are replaced with binary variables. Each $k_{i}$ requires $bl(f_{i}^{\text{max}})$ new binary variables, where $bl(x)$ denotes the bit-length of $x$.
  4. 4.
    In the final step, polynomial $F^{\prime}_{\text{Pen}}$ is determined according to Equation (13):
    $F^{\prime}_{\text{Pen}}=\sum_{i=0}^{m-1}(f^{\prime}_{i})^{2},$ (13)
    The standard components with a penalty are omitted because they are zero. Finally, the constant present in the polynomial is subtracted from the resulting polynomial.
The final matrix obtained has a size of $N=2728$ variables and $20598$ non-zero coefficients, which is $0.55\%$ of all matrix elements.
@@ 4.3 How to identify the problem and parameterize it?
Quantum service providers require data to identify the user. Among such data, we can distinguish:
@@
Using them, an untrusted provider can identify the client and the research area in which the client works. Identifying the area of study will allow the data obtained to be matched with one of the known problems. Problems vary significantly in size, density, and coefficient range. We assume that resources cannot freely change the connectivity (density) or size of the problem.
The lowest layer and critical element of the attack is the parameterization of the matrix of the selected optimization problem. The parameterization details will depend on the type of optimization task. However, the general principle remains the same. Below is an idea of how it can be realized for an algebraic attack on an $E_{0}$ cipher.
As described earlier, the selected optimization problem depends on the given keystream. To parameterize its matrix, we need to identify the coefficients of the matrix that depend on the $z_{i}$. As shown in Equation (12), only one of the equations depends on the bits of the keystream. We highlight two methods: the equation analysis dependent on $z_{i}$ and the algorithmic approach based on the methods of construction of the optimization problem, where an oracle $\phi$ is created.
The first method requires analyzing the equations. According to the method of transforming the algebraic attack described earlier, we focus on Equation (13). For each bit of the keystream, there will be a component in the final polynomial shown in:
$z_{t}+l_{t+1}+m_{t+7}+n_{t+1}+o_{t+7}+c^{0}_{t}-2K=0\Big{/}^{2}.$
After squaring, we get six coefficients depending on the keystream bits: $2z_{t}\cdot l_{t+1},2z_{t}\cdot m_{t+7},2z_{t}\cdot n_{t+1},2z_{t}\cdot o_{t+7},2z_{t}\cdot c_{t}^{0},-4z_{t}\cdot K$. Note that $K$ is an integer variable and should be replaced with a binary variable. The coefficients resulting from the squaring of other equations are constant. With the above knowledge and using Equation (6), equations dependent on the keystream occurring in the vector of biases can be computed.
The alternative method does not require direct analysis of equations. It is more generic and can be used universally for any problem for which the construction of the optimization problem is known. The coefficients of the matrix depend linearly on secret information. This is the situation for the $E_{0}$ cipher and most stream ciphers. Assume that an oracle $\phi$ is given which, for any keystream, will return an Ising model, as in Equation (15). The oracle can be constructed based on known publications on selected optimization problems, similar to the earlier description in Section 4.2. The chosen method requires $|z|+1$ queries to create a parameterized matrix for the keystream of length $|z|$.
To determine the coefficients depending on a specific bit of the keystream, we use all streams with a Hamming weight of 1. Such a stream can be denoted as $s_{i}$ and is presented as Equation (14):
@@
We denote the Ising models corresponding to the streams $s_{i}$ as $h_{i}$ and $J_{i}$. They are determined using the oracle $\phi$, as shown in Equation (15):
@@
Additionally, we denote $h_{\infty}$ and $J_{\infty}$ as the Ising model for a stream with Hamming weight equal to zero. Then, the parameterized problem is computed as follows:
$\begin{array}[]{c}h^{P}=h_{\infty}+\sum_{i=0}^{n}(h_{i}-h_{\infty})\cdot z_{i},\\ J^{P}=J_{\infty}+\sum_{i=0}^{n}(J_{i}-J_{\infty})\cdot z_{i},\end{array}$ (16)
where $z_{i}$ is a variable.
The idea of this parameterization method is illustrated in the example in Section 5.1.